OpenStudy (miihm):

How do you convert square feet to square yards? 24 Square feet to square yards. There are 9 square feet in 1 yard square.

OpenStudy (weasel123):

multiply the square feet by 9 i think

OpenStudy (miihm):

So 9 times nine? or 9 times 24? I tried nine times 24 and it said it was wrong and to try again.

OpenStudy (miihm):

I used the formula for conversion but it doesn't work with this problem. :(

OpenStudy (anonymous):

1 yard is 3 ft, yes?

OpenStudy (miihm):

No its says its nine.

OpenStudy (miihm):

Unless I have to use the actual metric sheet.

OpenStudy (anonymous):

The square root of 9 would be 3, right?

OpenStudy (miihm):

Yes.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 1~\color{grey}{\rm yard}=3~\color{grey}{\rm ft} }\) \(\color{#000000 }{ \displaystyle (1~\color{grey}{\rm yard})^2=(3~\color{grey}{\rm ft})^2 }\) \(\color{#000000 }{ \displaystyle 1~\color{grey}{\rm yard^2}=9~\color{grey}{\rm ft^2} }\) \(\color{#000000 }{ \displaystyle (1/9)~\color{grey}{\rm yard^2}=1~\color{grey}{\rm ft^2} }\)

OpenStudy (miihm):

Oh, that makes much more sense. Thank you.

OpenStudy (anonymous):

Aw...

OpenStudy (solomonzelman):

So, if you want to convert \(x\) square-feet to square-yards, then, you get. \(\color{#000000 }{ \displaystyle (1/9)~\color{grey}{\rm yard^2}=1~\color{grey}{\rm ft^2} }\) \(\color{#000000 }{ \displaystyle (x/9)~\color{grey}{\rm yard^2}=x~\color{grey}{\rm ft^2} }\)

OpenStudy (miihm):

I am still getting the answer wrong.

OpenStudy (miihm):

I must be forgetting somethings.

OpenStudy (anonymous):

All I could think is 3yd^2.

OpenStudy (miihm):

I finally figured it out. The answer was 24/9.

OpenStudy (solomonzelman):

yes, 24/9 square-yards.